OpenStudy (anonymous):

find the sum of the first 50 terms of the sequence.. a(n)=3n+2 do I just plug in 50? so the answer is 152?

OpenStudy (anonymous):

you need the sum 3(1) + 2 + 3(2) + 2 + 3(3) + 2

OpenStudy (anonymous):

no, that would be the 50th term, not the sum of all 50 terms

OpenStudy (anonymous):

You have to find the sum of all 50 elements {5, 8, 11, 14, 17, 20, 23, 26, 29, 32, 35, 38, 41, 44, 47, 50, 53, \ 56, 59, 62, 65, 68, 71, 74, 77, 80, 83, 86, 89, 92, 95, 98, 101, 104, \ 107, 110, 113, 116, 119, 122, 125, 128, 131, 134, 137, 140, 143, 146, \ 149, 152}

OpenStudy (anonymous):

Of course, you can add them term by term or you can think about a shorter way to do it

OpenStudy (anonymous):

i like your way better :)

OpenStudy (anonymous):

Think about \[ 1+2 +3 \cdots + (n-1)+n = \frac {n(n+1)}2 \]

OpenStudy (anonymous):

\[ \sum_{n=1}^{50} 3 n +2 = 2(50 )+ 3 \frac {50* 51}2=3925 \]

OpenStudy (anonymous):

@hi5tigtg58 , but how do you do this sum for the first 10,000 elements. You are not going to add them up term by term.

OpenStudy (mathmale):

Yes, and also about \[\sum_{1}^{n}1=n\] If \[a ^{n}=3n+2\] then the sum of the first n terms of this series could be expressed as \[\sum_{1}^{n}(3n+2)=3\sum_{1}^{n}n+2\sum_{1}^{n}1\]

OpenStudy (mathmale):

So, Hi-5, what would the value of n be in this case? How would you evaluate \[\sum_{1}^{n}n? \sum_{1}^{n}1?\]for that n?

OpenStudy (mathmale):

(That's two separate sums.)